# Efficient Parallel Algorithms for Inpainting-Based Representations of 4K Images

### v1

@@ 1.3 Paper Structure
We give a short review of homogeneous diffusion inpainting in Section 2. In Section 3, we present our Delaunay-based densification method for the spatial optimization. Section 4 then provides details on our tonal optimization methods as well as our strategy for constructing a good initial guess. We evaluate our spatial and tonal optimization methods experimentally in Section 5. Finally, we conclude our paper in Section 6 and discuss our ongoing work.
@@ 2 Homogeneous Diffusion Inpainting
In this section we briefly review homogeneous diffusion inpainting [14], which is our inpainting method of choice and is used for all results in this paper. Homogeneous diffusion inpainting is a simple and very popular method for inpainting-based compression. It is parameter-free, allows to establish a comprehensive data selection theory, and can lead to particularly fast algorithms. It has been shown that, if the inpainting data is carefully chosen, it can achieve very good results [31, 35, 43, 44]. For a more detailed review, see the first part of our companion paper [39].
@@ 2.1 Mathematical Formulation
Let us consider a continuous grayscale image $f:\Omega\to\mathbb{R}$, defined on a rectangular image domain $\Omega\subset\mathbb{R}^{2}$, which is known only on a small subset $K\subset\Omega$. Homogeneous diffusion inpainting aims at restoring $f$ in the inpainting domain $\Omega\setminus K$, by solving the Laplace equation on $\Omega\setminus K$ with reflecting boundary conditions on $\partial\Omega$, and with the known data $f|_{K}$ used as an interpolation constraint $u|_{K}=f|_{K}$:
$-\Delta u(\bm{x})$ $=0,$ $\bm{x}\in\Omega\setminus K,$ (1)
$u(\bm{x})$ $=f(\bm{x}),$ $\bm{x}\in K,$ (2)
$\partial_{\bm{n}}u(\bm{x})$ $=0$ $\bm{x}\in\partial\Omega,$ (3)
@@
where $\bm{I}\in\mathbb{R}^{N\times N}$ is the identity matrix, $\bm{C}:=\text{diag}(\bm{c})\in\mathbb{R}^{N\times N}$ is a diagonal matrix with the components of the inpainting mask on the diagonal, and $\bm{L}$ is the $5$-point stencil discretization of the negated Laplacian with reflecting boundary conditions. For an RGB color image the inpainting is performed on each of the three channels separately. Thus, we get three non-coupled linear systems of this type (the system matrix is the same for all of them).
@@ 3.1 Densification
[…]
@@
To start the densification, we need an initial mask $\bm{c}$ with a small amount of mask pixels in order to compute the first inpainting. We may obtain such an initial guess by using uniform random sampling or with another spatial optimization method, for example the Belhachmi approach [8]. In each iteration, we compute an inpainting $\bm{u}=\bm{A}^{-1}\bm{C}\bm{f}$ from the mask in the current iteration. We can then compute the error map $\bm{e}=|\bm{u}-\bm{f}|^{2}$, partition the domain using a Delaunay triangulation with the current mask pixels as vertices, and then integrate the error map within each triangle $\mathcal{T}_{i}$ to obtain the accumulated errors $e_{\mathcal{T}_{i}}$:
$e_{\mathcal{T}_{i}}=\sum_{j\in\mathcal{T}_{i}}e_{j}=\sum_{j\in\mathcal{T}_{i}}|u_{j}-f_{j}|^{2}.$ (5)
Since we need to introduce multiple pixels in a single iteration, we distribute them so that we insert at most one per triangle, and we choose the triangles in order of decreasing triangle error $e_{\mathcal{T}_{i}}$. The pixels are placed at the location of highest pointwise error within the triangles. This allows us to avoid clumping while at the same time having a scale-adaptive partition. We iterate this process until we reach the desired mask density. Algorithm 1 describes the Delaunay densification process, and Fig. 2 visualizes it.
@@
Input : Original image $\bm{f}$, number of iterations $n$, number of mask pixels added in first iteration $m_{0}$
@@
2:     Construct the Delaunay triangulation of the current mask pixels.
@@ 3.2 Implementation Details
Here we discuss some implementation components that can improve the mask construction process.
@@
For the initial mask in the first iteration, both [18] and [16] distribute the mask pixels uniformly at random. If we use only a very small amount of mask pixels in the initial mask, it does not have a substantial impact on the final densification result. However, as the initial mask density increases (i.e. as the number of densification steps decrease) the impact on the final result also increases. Generally, a uniform random mask is not optimal, as it is not adapted to image structures in any way. A much better distribution can be achieved by choosing the local mask density to be proportional to the Laplacian magnitude, as proposed by Belhachmi et al. [8]. They use a Floyd-Steinberg dithering [22] on the Laplacian magnitude of the reference image, however, such error diffusion algorithms are not well suited for GPU parallelization. Instead, we use a simple random dithering - we do a biased coin flip for every pixel to decide whether it should become a mask pixel - with pixel probabilities proportional to the Laplacian magnitude. While by itself this results in a worse reconstruction quality compared to Floyd-Steinberg dithering, the quality degradation in a Delaunay densification process is negligible, and the greater efficiency more than pays off for it.
@@
We compute the inpainting in each iteration with our multigrid ORAS method [39] which is currently the fastest homogeneous diffusion inpainting method for highly parallel architectures such as GPUs. This allows us to perform up to $20$ iterations in under $0.5$ seconds with Delaunay densification.
@@ 4 Tonal Optimization with Domain Decomposition
$\underset{\bm{g}\in\mathbb{R}^{N}}{\operatorname*{argmin}}\|\bm{f}-\bm{B}\,\bm{g}\|_{2}^{2}.$ (6)
Since grayscale or color values $\bm{g}_{i}$ at pixels $i\in\Omega\setminus K$ outside the inpainting mask $\bm{c}$ have no influence on the inpainting result we simply set them to 00. Let us now discuss, how we solve the resulting least squares problem efficiently.
[…]

### v2

@@ 1.3 Paper Structure
We give a short review of homogeneous diffusion inpainting in Section 2 and in Section 3 we give an overview of the Multigrid ORAS inpainting solver that we use in our spatial and tonal optimization methods. In Section 4, we present our Delaunay-based densification method for the spatial optimization. Section 5 provides details on our tonal optimization methods as well as our strategy for constructing a good initial guess. We evaluate our spatial and tonal optimization methods experimentally in Section 6. Finally, we conclude our paper in Section 7 and discuss our ongoing work.
@@ 2 Homogeneous Diffusion Inpainting
In this section we briefly review homogeneous diffusion inpainting [19], which is our inpainting method of choice and is used for all results in this paper. Homogeneous diffusion inpainting is a simple and very popular method for inpainting-based compression. It is parameter-free, allows to establish a comprehensive data selection theory, and can lead to particularly fast algorithms. It has been shown that, if the inpainting data is carefully chosen, it can achieve very good results [42, 46, 53, 54].
@@ 2.1 Mathematical Formulation
Let us consider a continuous grayscale image $f:\Omega\to\mathbb{R}$, defined on a rectangular image domain $\Omega\subset\mathbb{R}^{2}$. We store only a fraction of the full image data on a small subset $K\subset\Omega$. Homogeneous diffusion inpainting aims at restoring $f$ in the inpainting domain $\Omega\setminus K$, by solving the Laplace equation on $\Omega\setminus K$ with reflecting boundary conditions on $\partial\Omega$, and with the known data $f|_{K}$ used as an interpolation constraint $u|_{K}=f|_{K}$:
(1) $-\Delta u(\bm{x})$ $=0,$ $\bm{x}\in\Omega\setminus K,$
(2) $u(\bm{x})$ $=f(\bm{x}),$ $\bm{x}\in K,$
(3) $\partial_{\bm{n}}u(\bm{x})$ $=0$ $\bm{x}\in\partial\Omega,$
@@
where $\bm{I}\in\mathbb{R}^{N\times N}$ is the identity matrix, $\bm{C}:=\text{diag}(\bm{c})\in\mathbb{R}^{N\times N}$ is a diagonal matrix with the components of the inpainting mask on the diagonal, and $\bm{L}$ is the $5$-point stencil discretization of the negated Laplacian with reflecting boundary conditions. If the inpainting mask is non-empty, i.e. $\bm{c}\neq\bm{0}$, the inpainting problem has a unique solution since $\bm{A}$ is non-singular in that case (for more details on the matrix structure see [53]). For an RGB color image the inpainting is performed on each of the three channels separately. Thus, we get three non-coupled linear systems of this type (the system matrix is the same for all of them).
Homogeneous Diffusion Inpainting and Image Edges
Edges are important semantic and mathematical features in images. Thus, a simple and appealing idea is to reconstruct an image from its edge information. When the Laplacian $\bm{L}\bm{f}$ is considered as an edge detector, this idea is realized as homogeneous diffusion inpainting [19]. If $\bm{L}\bm{f}$ is available exactly, the image can be recovered up to mean gray value as $\bm{L}^{+}(\bm{L}\bm{f})$, where $\bm{L}^{+}$ is the Moore-Penrose inverse. Since $\bm{L}$ has a high-pass profile, $\bm{L}^{+}$ has a low-pass character and (4) fills in information smoothly between mask points.
In practice, we do not want to store $\bm{L}\bm{f}$ exactly, leading to the sparse approximation problem with the inpainting from (4). Earlier approaches [1, 7, 19, 25, 34, 52, 86] for homogeneous diffusion inpainting used the pixels left and right of the edges as known data, in order to preserve them in the reconstruction. While modern approaches usually do not directly use the edges to determine the inpainting data, the optimized inpainting masks still cluster around prominent image edges. For an illustration of this effect, see Figure 12. Further theoretical justifications for choosing reconstruction data based on the Laplacian can be found in the work of Belhachmi et al. [8].
We also note that the choice of reflecting boundary conditions (3) is standard in image processing, as it produces the best results for the compression of natural images. In fact the matrix $\bm{L}$ has eigenvectors the rows of the DCT-II transform [75], which is known to have the best energy compaction properties for natural images, out of all the DFT/DST/DCT variants.
Alternative Inpainting Operators
An alternative class of sparse inpainting operators use locally supported positive definite kernels [23, 60, 82]. While the latter has considerable advantages since it does not have to deal with global interactions, there are non-trivial challenges associated with this approach. The kernels come with various parameters that have to be optimized per mask point. This necessitates multiple iterations to produce a single inpainting [23], that is neither over-blurred – due to the kernel supports being too large, nor has holes – due to the kernel supports being too small.
Other options include nonlinear inpainting methods such as anisotropic diffusion inpainting [81]. Such methods can achieve state-of-the-art quality [32], however, that comes at a considerable computational cost. We have left the study of these alternatives as future work, and have focused instead on constructing the first complete framework for data optimization applicable to sparse 4K image inpainting.
3 Multigrid ORAS Inpainting Solver
To keep our manuscript self-contained, we describe the Multigrid ORAS inpainting solver that we use for the inpainting problems that occur in our proposed data optimization methods. It consists of a domain decomposition method embedded into a multigrid framework. For more details on the Multigrid ORAS inpainting solver, we refer the reader to [49].
3.1 Optimized Restricted Additive Schwarz Method
Figure 1: Domain Decomposition Example. The domain is divided into four overlapping subdomains. The subdomain $\Omega_{1}$ is highlighted in blue.
The optimized restricted additive Schwarz (ORAS) method [74], in its discrete formulation, is an iterative technique for solving linear systems that arise from the discretization of boundary value problems, such as Eq. 4. It is one of the simplest domain decomposition methods [28, 78] and is well suited for parallelization. As a first step, the image domain $\Omega$ with $N$ pixels is divided into $k\in\mathbb{N}$ overlapping subdomains $\Omega_{1},...,\Omega_{k}\subset\Omega$, such that $\cup_{i=1}^{k}\Omega_{i}=\Omega$. An example of such an overlapping division into four subdomains is shown in Fig. 1.
Starting with an initial approximation $\bm{u}^{0}\in\mathbb{R}^{N}$ that fulfills the interpolation condition $\bm{C}\bm{u}^{0}=\bm{C}\bm{f}$ at the known locations, in each iteration $n$, we try to correct the current approximation $\bm{u}^{n}$ by locally computing correction terms for every subdomain. We compute the corrections by solving local variants of the global linear system
(5) $\bm{A}_{i}\bm{v}_{i}^{n}=\bm{R}_{i}\bm{r}^{n},$
where the right-hand side consists of the global residual $\bm{r}^{n}=\bm{b}-\bm{A}\bm{u}^{n}$ restricted to the subdomains. We use matrices $\bm{R}_{i}\in\mathbb{R}^{|\Omega_{i}|\times N}$ to restrict global vectors to the local domains by extracting the values from the subdomain $\Omega_{i}$ and ignoring all values outside of it. The local system matrices $\bm{A}_{i}$ are given by restricting the global system matrix $\bm{A}$ to the subdomains $\Omega_{i}$ with additional boundary conditions at the subdomain boundaries which are not part of the global image boundary. At these so-called inner subdomain boundaries, we impose mixed Robin boundary conditions, which improve the convergence speed compared to simpler Dirichlet boundary conditions [74].
The next iterate $\bm{u}_{n+1}$ is given by extending the local corrections $\bm{v}_{i}$ from the subdomain $\Omega_{i}$ to the global domain $\Omega$ and adding them to the old iterate $\bm{u}_{n}$:
(6) $\bm{u}^{n+1}=\bm{u}^{n}+\sum_{i=1}^{k}\bm{R}_{i}^{\top}\bm{D_{i}}\bm{v}_{i}^{n}$
For the extension, we use the transposes of the restriction matrices $\bm{R}^{\top}_{i}$. However, in order to achieve convergence, we also have to weight the local corrections in the overlapping regions before adding them to the global solution [29, 33], to compensate for the fact that we add multiple corrections in the overlapping regions. To this end we define diagonal nonnegative weight matrices $\bm{D}_{i}\in\mathbb{R}^{\lvert\Omega_{i}\rvert\times\lvert\Omega_{i}\rvert}$, such that
(7) $\bm{I}=\sum_{i=1}^{k}\bm{R}_{i}^{\top}\bm{D}_{i}\,\bm{R}_{i},$
where $\bm{I}\in\mathbb{R}^{N\times N}$ is the identity matrix of size $N\times N$.
3.2 Embedding ORAS in Multigrid
Since the ORAS method only allows communication between neighboring blocks, many iterations are necessary to spread information over large distances. This means that high frequency components of the error are reduced much faster than low frequency components, which leads to an overall slow convergence rate. As a remedy, domain decomposition methods are usually used together with a two-level scheme that computes a correction on a very coarse grid to couple all subdomains together [28, 78]. In our case, where the influence area of each mask pixel is already approximately localized, such a global coupling would be to coarse and result in a poor performance. Instead, the more fine-grained coupling of multigrid methods [14, 15, 17, 26, 36, 79, 83] can be more beneficial. Multigrid methods transfer the problem to coarser grids where the low frequencies appear as higher frequencies, which can be reduced more efficiently with iterative solvers such as Jacobi [69], Gauss-Seidel [69], or the ORAS method [74].
Two-Grid Cycle
We first give an overview over the simple two-grid cycle, which we then extend to use multiple resolution levels. Our two grids are the original fine grid with grid spacing $h$, and a coarser grid with spacing $H=2h$.
We start the cycle $k+1$ with an approximation $\bm{u}_{h}^{k}$ of the solution $\bm{u}_{h}$ of the problem $\bm{A}_{h}\bm{u}_{h}=\bm{b}_{h}$, where $\bm{A}_{h}$ is the original system matrix and $\bm{b}_{h}$ the right-hand side on the original grid. The goal of the two-grid cycle is to find a good approximation of the true error $\bm{u}_{h}-\bm{u}_{h}^{k}$, by decimating different frequencies of the error over different grids. We reduce the high frequencies by performing a few iterations $\vartheta_{\text{pre}}$ with a smoother: an iterative solver that dampens the high frequencies efficiently - e.g. damped Jacobi, Gauss-Seidel, or in our case ORAS. This results in the approximation of the solution
(8) $\bm{u}_{h}^{k+1/3}=\text{pre-smooth}(\bm{A}_{h},\bm{b}_{h},\bm{u}_{h}^{k},\vartheta_{\text{pre}}).$
With enough iterations $\vartheta_{\text{pre}}$, the error will have only negligible high-frequency components and we can transfer it almost perfectly to the coarser grid with a restriction matrix $\bm{R}^{H}_{h}$. While we do not know the true error in practice, we know the residual and its relation to the error:
(9) $\bm{r}^{k}_{h}=\bm{b}_{h}-\bm{A}_{h}\bm{u}_{h}^{k+1/3}=\bm{A}_{h}\bm{u}_{h}-\bm{A}_{h}\bm{u}_{h}^{k+1/3}=\bm{A}_{h}\left(\bm{u}_{h}-\bm{u}_{h}^{k+1/3}\right)=\bm{A}_{h}\bm{e}_{h}^{k}.$
Instead of solving this equation on the fine grid, we transfer it to the coarser grid and solve it for the coarse error $\bm{e}_{H}^{k}$ as
(10) $\bm{A}_{H}\bm{e}_{H}^{k}=\bm{r}^{k}_{H}.$
Here $\bm{A}_{H}$ is an analogue of $\bm{A}_{h}$, obtained by discretizing the inpainting problem on the coarser grid, and $\bm{r}^{k}_{H}=\bm{R}^{H}_{h}\bm{r}^{k}_{h}$ is the coarse residual, obtained by restricting the original residual to the coarser grid. Since the matrix $\bm{A}_{H}$ is twice smaller in each dimension compared to $\bm{A}_{h}$, the computational cost is also reduced. We can then transfer $\bm{e}_{H}^{k}$ to the fine grid by using a prolongation matrix $\tilde{\bm{e}}^{k}_{h}=\bm{P}^{h}_{H}\bm{e}^{k}_{H}$ and use it to correct our approximation of the solution:
(11) $\bm{u}_{h}^{k+2/3}=\bm{u}^{k+1/3}+\tilde{\bm{e}}^{k}_{h}=\bm{u}^{k+1/3}+\bm{P}^{h}_{H}\bm{e}^{k}_{H}.$
A subsequent post-smoothing with $\vartheta_{\text{post}}$ iterations is applied to smooth high frequency error components potentially introduced by the correction step:
(12) $\bm{u}^{k+1}_{h}=\text{post-smooth}(\bm{A}_{h},\bm{b}_{h},\bm{u}_{h}^{k+2/3},\vartheta_{\text{post}}).$
Extension to More Resolution Levels
We extend the two-grid cycle to multiple levels, which allows to efficiently dampen different frequencies of the error at different resolutions. The solve on the coarse level is replaced by another correction step on an even coarser level. We iterate this process recursively until the desired number of resolution levels are reached. This type of hierarchical application of the two-grid cycle is termed a V-cycle.
Additionally we speed up the computation even further by providing a good initialization. The latter is obtained by computing an inpainting solution on a coarser grid and prolongating it to the fine resolution. We extend this to to a complete coarse-to-fine initialization strategy. While this is usually used with a complete V-cycle on each resolution level, we instead skip the V-cycles on the lower levels and just use a single pre-smoothing iteration in each lower level. The resulting reduced full multigrid (FMG) scheme, is depicted in Fig. 2. It decreases the amount of costly restriction and prolongation operations, while still offering sufficient convergence for homogeneous diffusion inpainting.
Figure 2: Reduced Full Multigrid Scheme. The initial guess is constructed in a coarse-to-fine manner, also known as one-way or cascadic multigrid [14]. Then we continue with additional V-cycle correction steps (a single V-cycle is visualized above).
3.3 Implementation Details
In our ORAS method we decompose the image domain into multiple overlapping blocks of size $32\times 32$ pixels with an overlap of $6$ pixels. These values are optimized for our Nvidia GeForce GTX 1080 Ti GPU, which results in $36852$ local problems in each iteration of the ORAS method for a 4K color image on the finest resolution layer. The local problems are solved using a simple conjugate gradient algorithm [69]. We stop the local CG iterations on a block when the squared 2-norm of the residual has been reduced to a fixed fraction of the squared global residual norm. Compared to a stopping criterion based on the local relative residual norm, this leads to a more uniform convergence, since all blocks have the same squared residual norm after each ORAS iteration. This also avoids using unnecessarily many CG iterations in blocks that are already sufficiently converged.
For our multigrid embedding, we use a single V-cycle on the finest level with a single ORAS iteration as pre- and post-smoothing. The number of resolution levels is chosen, such that the coarsest level consists of a single ORAS block. For the restriction of the inpainting mask we consider a coarse pixel to be a known pixel, if at least one of its four corresponding fine resolution pixels is a known pixel. For each coarse mask pixel its value is obtained as an average over the corresponding fine resolution pixel values. This method offers the advantage of increasing the mask density with each level, which in turn leads to faster convergence at the coarser levels.
@@ 4.1 Densification
(a) original image trui
(b) Voronoi tessellation
(c) Delaunay triangulation
@@
To start the densification, we need a non-empty initial mask $\bm{c}$ with a small amount of mask pixels in order to compute the first inpainting. We may obtain such an initial guess by using uniform random sampling or with another spatial optimization method, for example the Belhachmi approach [8]. In each iteration, we compute an inpainting $\bm{u}=\bm{A}^{-1}\bm{C}\bm{f}$ from the mask in the current iteration. We can then compute the error map $\bm{e}=|\bm{u}-\bm{f}|^{2}$, partition the domain using a Delaunay triangulation with the current mask pixels as vertices, and then integrate the error map within each triangle $\mathcal{T}_{i}$ to obtain the accumulated errors $e_{\mathcal{T}_{i}}$:
(13) $e_{\mathcal{T}_{i}}=\sum_{j\in\mathcal{T}_{i}}e_{j}=\sum_{j\in\mathcal{T}_{i}}|u_{j}-f_{j}|^{2}.$
Since we need to introduce multiple pixels in a single iteration, we distribute them so that we insert at most one per triangle, and we choose the triangles in order of decreasing triangle error $e_{\mathcal{T}_{i}}$. The pixels are placed at the location of highest pointwise error within the triangles. This allows us to avoid clumping while at the same time having a scale-adaptive partition. We iterate this process until we reach the desired mask density. Algorithm 1 describes the Delaunay densification process, and Fig. 4 illustrates it.
Comparison to Spatial Optimization for Delaunay Based Inpainting
Various methods rely on (at a first glance) similar strategies for greedily optimizing the inpainting mask. Specifically in the context of the used Delaunay partition, the work of Demaret et al. [24], Adams [3], and Marwood et al. [55] use linear splines over Delaunay triangulations for sparse image reconstruction. Their methods are specifically adapted to the linear splines Delaunay inpainting that is motivated from the desire of not storing the connectivity of a triangulation, while our method could easily be generalized for other inpainting methods. Our method is also more sophisticated, as we integrate the errors over the adaptive partition as a predictor for the potential incurred error from removing or adding a mask pixel. Last but not least, our method is also significantly faster, as their method requires approximately $4$ minutes for $23,092$ initial data points, which is about $3$ times lower than for a $256\times 256$ image.
@@
Input : Original image $\bm{f}$, number of iterations $n$, number of mask pixels added in first iteration $m_{0}$, factor to increase the number of mask-pixels added per iteration $t$
@@
2:     Construct the Delaunay triangulation $\{\mathcal{T}_{j}\}$ of the current mask pixels.
@@ 4.2 Implementation Details
Here we discuss implementation details that improve the mask construction process from both a quality and runtime perspective.
@@
For the initial mask in the first iteration, both [23] and [21] distribute the mask pixels uniformly at random. If we use only a very small amount of mask pixels in the initial mask, it does not have a substantial impact on the final densification result. However, as the initial mask density increases (i.e. as the number of densification steps decrease) the impact on the final result also increases. Generally, a uniform random mask is not optimal, as it is not adapted to image structures in any way. A much better distribution can be achieved by choosing the local mask density to be proportional to the Laplacian magnitude, as proposed by Belhachmi et al. [8]. They use a Floyd-Steinberg dithering [31] on the Laplacian magnitude of the reference image, however, such error diffusion algorithms are not well suited for GPU parallelization. Instead, we use a simple random dithering on the Laplacian magnitude - we do a biased coin flip for every pixel to decide whether it should become a mask pixel - with pixel probabilities proportional to the Laplacian magnitude. While by itself this results in a worse reconstruction quality compared to Floyd-Steinberg dithering, the quality degradation in a Delaunay densification process is negligible, and the greater efficiency more than pays off for it.
@@
By using our multigrid ORAS solver to compute the inpainting in each iteration, we are able to perform up to $20$ iterations in under $0.5$ seconds with our Delaunay densification.
@@ 5 Tonal Optimization with Domain Decomposition
(14) $\underset{\bm{g}\in\mathbb{R}^{N}}{\operatorname*{argmin}}\|\bm{f}-\bm{B}\,\bm{g}\|_{2}^{2}.$
Since grayscale or color values $\bm{g}_{i}$ at pixels $i\in\Omega\setminus K$ outside the inpainting mask $\bm{c}$ have no influence on the inpainting result we simply set them to 00. The choice of the $2$-norm is especially appealing here (and rather standard in image processing) as it results in a linear least squares problem. Other options that result in linear least squares would include weighted $2$-norms $\|\cdot\|_{\bm{G}}$, however, in our case we want the reconstructed image $\bm{B}\bm{g}$ to look as close as possible to the original $\bm{f}$ overall, so introducing a weighting matrix does not provide any benefits. A weighted norm could be used potentially in a foveated rendering setting.
In the following subsection we discuss how we solve the resulting least squares problem efficiently.
(a) original image trui
(b) no tonal opt.: 141.48 MSE
(c) tonal opt.: 63.78 MSE